Obtain the Maclaurin series for f(x)=(1/(2-x))
1
Expert's answer
2020-05-18T19:46:51-0400

The maclaurin series is given by:

f(x)= f(0) + f'(0)x + "\frac {x^2 f''(0)}{2!}" + "\frac {x^3 f'''(0)}{3!}" + ..... +"\frac {x^n f^n(0)}{n!}+....."


f = (2-x)-1

f' = (2-x)-2

f''' = -2(2-x)-3

f'''' = 6(2-x)-4

........

f(0) = 1/2

f'(0) = -24

f''(0) = 6.2-4

......

maclaurin series = "\frac {1}{2}" + "\frac {x}{4} - \frac{x^2}{2^3} + \frac{x^3}{2^4} - \frac{x^4}{2^5}+ ....."